7. An urn contains four black and eight white balls. A sample of five balls is selected from the urn repeatedly, each time with replacement of the selected balls (thus restoring the urn to its original state). Let E be the event that the selected sample contains at most two white balls. (a) Find the probability of E (b) What is the expected number of selections until E happens? Name (c) What is the probability that E happens for the first time on the 10th (d) What is the probability that B happens the tenth time on the 6oth (e) What is the expected number of selections until E happens for the the random variable used. selection? selection? tenth time? Name the random variable used.
Answer #1

a) here as probability of a white ball =8/12=2/3

therefore from binomial distribution P(X<=2)=P(X=0)+P(X=1)+P(X=2)

=5C0(2/3)0(1/3)5+5C1(2/3)1(1/3)4+5C2(2/3)2(1/3)3=0.209877

b)

this follows geometric distribution with parameter p=0.209877

expected selections =1/p=1/0.209877=4.76

c)

P(first time on 10th selection)=P(not happened for first 9 time and happens on 10th time)=(1-0.209877)9*(0.209877)

=0.025189

d)

Probability=P(E happens exactly 9 times in first 59 selection and happens on 60th selection)

=59C9(0.209877)9*(1-0.209877)=0.01598

e)

this follows negative binomial with paramter n=10 and p=0.209877

expected number =r/p=10/0.209877=47.64696